Jason invested a total of 39000 into two accounts, one yielding 5% annual interest and the other yielding 10% annual interest (a more risky investment). If he received a total of 3025 interest at the end of the year, how much did he invest in each account
Answer by fractalier(6550)

Call the amount invested at 5%, x. Then the amount invested at 10% would be 39000-x. The setup looks like this
.05x + .10(39000-x) = 3025
.05x + 3900 - .1x = 3025
-.05x = -875
x = 875/.05 = $17,500 at 5%
and then
39000-x = $21,500 at 10%